Database last updated on Mon, 02 Feb 2026 01:10:59 +0000 by localhost

Instrument constant changes for the week before 2023344 (3 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 15 rows)

Measurements and tests summary.


_jday_loc_serial
344TORONTO#254
(0 sec)

Number of days since last dome/window cleaning (0 sec, 1 rows)

Missing files (0 sec, 1 rows)

Seriallast datadays ago
25420233450-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
254501000000

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (1 sec)

No ozone measurements to display.

Unsorted comments (0 sec, 9906 rows)

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk